# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

### v1

@@ IV-A3 Discussion
Compared to existing purely sequential insertion procedures [15, 21, 22], PSSI can typically obtain higher-quality feasible routes with reasonable additional computational cost, as demonstrated in the experiments (see Section V-D). The applicability of PSSI extends beyond EVRP-TW-SPD to various EVRP variants sharing certain features: i) partial recharges, ii) customers with time windows, and iii) a homogeneous network of charging stations. For instance, PSSI could be applied to scenarios involving multiple depots, heterogeneous vehicle fleets, or even dynamic routing where customer demands or traffic conditions change in real-time.
@@ IV-B Cross-Domain Neighborhood Search (CDNS)
We propose cross-domain neighborhood search (CDNS), a novel local search procedure that explores two distinct solution spaces simultaneously: the solution space of EVRP-TW-SPD and the solution space of its non-electric counterpart VRP-TW-SPD [19], where some solutions may be infeasible for EVRP-TW-SPD. This cross-domain exploration offers two key advantages. First, by expanding beyond the feasible solution space of EVRP-TW-SPD, CDNS potentially discovers higher-quality solutions that might be overlooked by conventional single-domain local search. Second, exploring the solution space of VRP-TW-SPD enables the efficient reuse of well-established local search procedures for VRP-TW-SPD. We will further discuss the broader implications of this cross-domain approach beyond EVRP-TW-SPD at the end of Section IV-B.
@@ 15       $\bm{S}\leftarrow$ find the best-improving solution in the EVRP-TW-SPD neighborhood of $\bm{S}$;
Algorithm 2 outlines the framework of CDNS. It consists of two sub-procedures. The first, called aggressive local search (ALS), explores the solution space of VRP-TW-SPD. As its name suggests, it temporarily ignores the electricity constraint (5g) and repeatedly applies existing local search procedures for VRP-TW-SPD to find the solution with maximum improvement (i.e., the best-improving solution) in the neighborhood of the current solution (lines 2-2). When ALS can no longer improve the solution, CDNS switches to conservative local search (CLS) which searches for improvements within the solution space of EVRP-TW-SPD (lines 2-2). Below we detail these two sub-procedures.
@@ IV-B1 Aggressive Local Search (ALS)
When a feasible solution $\bm{S}$ to EVRP-TW-SPD enters ALS (lines 2-2), the first step is to remove all charging stations in $\bm{S}$, resulting in a feasible solution $\bar{\bm{S}}$ to VRP-TW-SPD (line 2). Then, an existing local search procedure for VRP-TW-SPD is repeatedly applied to find the best-improving solution in the neighborhood of $\bar{\bm{S}}$ (line 2). Here, we employ the local search procedure of MATE [19], one of the state-of-the-art algorithms for solving VRP-TW-SPD. This procedure involves multiple move operators and constant-time move evaluation (see the paper of MATE for details). Then, for each found best-improving solution, PSSI is first applied to transform it into a feasible EVRP-TW-SPD solution (lines 2-2). If this transformation is not possible, then ALS would terminate (i.e., the second termination condition in line 2). If successful, the resulting feasible solution would replace the current solution $\bm{S}$ when it has better quality (line 2). ALS would terminate when no further improvement is possible (i.e., the first termination condition in line 2).
@@ IV-B2 Conservative Local Search (CLS)
Unlike ALS, CLS performs local search directly in the solution space of EVRP-TW-SPD. It repeatedly identifies the best-improving solution in the current solution’s neighborhood until no further improvement is possible (lines 2-2). The neighborhood is defined by the following five move operators.
@@ IV-B3 Discussion
By exploring two distinct solution spaces, CNDS typically finds higher-quality solutions than using either solution space alone, as demonstrated in our experiments (see Section V-D). The idea of CDNS can be readily generalized to other EVRP variants by simultaneously exploring both electric and non-electric neighborhoods. This dual-domain approach is particularly valuable since non-electric variants are typically more extensively studied in the literature, providing a rich pool of well-established local search procedures that can be adapted for solving various EVRPs efficiently.
@@ 6       $\bm{S}^{\prime}\leftarrow\text{CDNS}(\bm{S}^{\prime})$;
9      if $\bm{S}$ is not improved in the last $G$ iterations then
@@ 21                   $\bm{S}^{\prime}\leftarrow\text{Select\_Best}(\bm{S}^{\prime},\bm{S}_{1},\ldots,\bm{S}_{N})$;
23            until $\bm{S}^{\prime}$ is not improved in the last $G$ generations;
@@
After initialization, HMA applies the destroy-repair operator as described in [19] (a large-step perturbation) on the individual solution $\bm{S}$ (line 3). This destroy-repair operator involves removing $\rho$ correlated customers from the solution, where the random number $\rho\in[\omega_{1}|\bm{V}|,\omega_{2}|\bm{V}|]$, and $\omega_{1}$, $\omega_{2}$ are parameters with $10<\omega_{1}\leq\omega_{2}<$ 1. The operator then reinserts these customers using a regret-based heuristic that evaluates the expected cost of inserting a node in a future iteration rather than in the current one to assess the quality of potential insertions. The process is followed by CDNS, ensuring that a new local optimum is reached to potentially improve $\bm{S}$ (lines 3-3). If $\bm{S}$ does not show improvement over a specified number of consecutive iterations, denoted as $G$, HMA initiates the population-based MS (lines 3-3).
@@
After initialization, MS enters an evolutionary process (lines 3-3). In each generation (lines 3-3), each individual in the population is selected once as parent $\bm{p}_{1}$ and once as parent $\bm{p}_{2}$, in a random order (lines 3 and 3). For each pair of parent solutions, MS generates an offspring using the route-assembly-regret-insertion (RARI) crossover [19] (line 3), and then improves it through CDNS (line 3). After that, the better one among the offspring and $\bm{p}_{1}$ replaces the individual selected as $\bm{p}_{1}$ (line 3). After MS, a new candidate solution $\bm{S}^{\prime}$ is obtained with the aim of improving $\bm{S}$ (lines 3-3). If the quality of $\bm{S}$ is still not improved by MS, then HMA would output $\bm{S}$ and terminate (line 3); otherwise, HMA proceeds to the next iteration of LNS (line 3).
When presented with a large-scale problem instance (with more than 200 customers) to solve, we make two adjustments to HMA to improve its efficiency. First, only ALS is used in CDNS. Second, problem decomposition is employed in MS. Specifically, the decomposition breaks down the main problem $\bm{G}=(\bm{V}^{\prime},\bm{E})$ into $k$ smaller subproblems $\bm{G}_{1},\ldots,\bm{G}_{k}$ when entering MS (line 3). A subproblem $\bm{G}_{i}=(\bm{V}_{i}^{\prime},\bm{E}_{i})$ with $\bm{V}_{i}^{\prime}=\{0\}\cup\bm{V}_{i}\cup\bm{F}$, $\bm{E}_{i}=\{(p,q)|p,q\in\bm{V}_{i}^{\prime},p\neq q\}$, where $\bm{V}=\bigcup_{i=1}^{k}\bm{V}_{i}\text{ and }\bm{V}_{i}\cap\bm{V}_{j}=\emptyset\ \forall i\neq j,\ 1\leq i,j\leq k$. Here, we use the barycenter clustering decomposition (BCD) [33] to determine the partition of $\bm{V}_{i}$. The barycenter of a route $\bm{R}=(n_{1},n_{2},\ldots,n_{L})$ is defined as $(\overline{x},\overline{y})=\frac{1}{L-1}\sum_{i=1}^{L-1}(x_{i},y_{i}),$ where $(x_{i},y_{i})$ is the Cartesian plane coordinate of node $n_{i}$. For each subproblem $\bm{G}_{i}$, the same MS is applied (lines 3-3). After all MS procedures are finished, the solutions from all subproblems are assembled to construct a complete solution $\bm{S}^{\prime}$ for the original problem.
@@
Parameter $G$ $N$ $\alpha$ $B$ $sr$ $w_{1},w_{2}$
$akb_{small}$ 20 9 3 5 $1.0$ $0.2,0.4$
$akb_{medium}$ 20 4 3 5 $0.5$ $0.1,0.2$
$jd$ 20 4 3 5 $0.1$ $0.05,0.05$
@@ V Experiments
To evaluate the effectiveness of HMA, we compared it with the recent algorithms for solving EVRP-TW-SPD, Adapt-CMSA-STD [11] and Adapt-CMSA-SETCOV [17], using the existing benchmark set. Then, we introduced a new benchmark set derived from a real-world application that includes large-scale instances. The evaluation results of HMA on the new benchmark were also reported. Finally, we conducted a comprehensive ablation study to assess the effectiveness of PSSI, CDNS, and the hybrid search framework. The source code of HMA and the new benchmark set are open-sourced anonymously at https://anonymous.4open.science/r/EVRP-TW-SPD-HMA-code-dataset.
@@
Instance m best time m best avg time m best avg time m best avg time
c101C5 2 2257.75 0.43 2 2257.75 2257.75 0.02 2 2257.75 2257.75 0.01 2 2257.75 2257.75 0.38
c103C5 1 1175.37 0.41 1 1175.37 1175.37 0.54 1 1175.37 1175.37 0.68 1 1175.37 1175.37 0.21
c206C5 1 1242.56 0.58 1 1242.56 1242.56 0.03 1 1242.56 1242.56 0.00 1 1242.56 1242.56 0.19
c208C5 1 1158.48 0.08 1 1158.48 1158.48 0.01 1 1158.48 1158.48 0.00 1 1158.48 1158.48 0.24
r104C5 2 2136.69 0.02 2 2136.69 2136.69 0.07 2 2136.69 2136.69 0.01 2 2136.69 2136.69 0.93
r105C5 2 2156.08 0.03 2 2156.08 2156.08 0.01 2 2156.08 2156.08 0.00 2 2156.08 2156.08 1.28
r202C5 1 1128.78 0.06 1 1128.78 1128.78 9.00 1 1128.78 1128.78 0.00 1 1128.78 1128.78 0.24
r203C5 1 1179.06 0.03 1 1179.06 1179.06 0.22 1 1179.06 1179.06 0.05 1 1179.06 1179.06 0.17
rc105C5 2 2233.77 2.18 2 2233.77 2233.77 0.09 2 2233.77 2233.77 0.04 2 2233.77 2233.77 0.86
rc108C5 2 2253.93 0.19 2 2253.93 2253.93 0.01 2 2253.93 2253.93 0.00 2 2253.93 2253.93 0.58
rc204C5 1 1176.39 0.25 1 1176.39 1176.39 0.07 1 1176.39 1176.39 0.01 1 1176.39 1176.39 0.21
rc208C5 1 1167.98 0.12 1 1167.98 1167.98 0.52 1 1167.98 1167.98 0.03 1 1167.98 1167.98 0.19
c101C10 3 3388.25 76.74 3 3388.25 3388.25 0.28 3 3388.25 3388.55 0.33 3 3388.25 3388.25 1.70
c104C10 2 2273.93 2.16 2 2273.93 2273.93 0.48 2 2273.93 2273.93 29.02 2 2273.93 2273.93 1.10
c202C10 1 1304.06 6.34 1 1304.06 1304.06 0.45 1 1304.06 1304.06 0.10 1 1304.06 1304.06 1.40
c205C10 2 2228.28 0.08 2 2228.28 2228.28 10.82 2 2228.28 2228.28 0.03 2 2228.28 2228.28 0.45
r102C10 3 3249.19 0.56 3 3249.19 3249.19 18.22 3 3249.19 3249.19 0.01 3 3249.19 3249.19 1.03
r103C10 2 2206.12 17.49 2 2206.12 2206.12 5.27 2 2206.12 2206.12 22.01 2 2206.12 2206.12 3.38
r201C10 1 1241.51 61.20 1 1241.51 1241.51 22.71 1 1241.51 1241.51 15.97 1 1241.51 1241.51 3.29
r203C10 1 1218.21 1.57 1 1218.21 1218.21 14.06 1 1218.21 1218.21 27.17 1 1218.21 1218.21 0.97
rc102C10 4 4423.51 0.15 4 4423.51 4423.51 0.24 4 4423.51 4423.51 0.01 4 4423.51 4423.51 3.27
rc108C10 3 3345.93 3.72 3 3345.93 3345.93 14.61 3 3345.93 3345.93 0.01 3 3345.93 3345.93 1.29
rc201C10 1 1412.86 22756.49 1 1412.86 1412.86 1.97 1 1412.86 1412.86 0.93 1 1412.86 1412.86 6.03
rc205C10 2 2325.98 0.13 2 2325.98 2325.98 0.13 2 2325.98 2325.98 0.45 2 2325.98 2325.98 0.74
c103C15 3 3348.46 5046.46 3 3348.46 3348.47 48.19 3 3348.46 3348.46 3.76 3 3348.46 3348.46 3.30
c106C15 3 3275.13 0.90 3 3275.13 3275.13 4.56 3 3275.13 3275.13 32.34 3 3275.13 3275.13 2.42
c202C15 2 2383.62 43.74 2 2383.62 2383.62 12.73 2 2383.62 2393.39 5.27 2 2383.62 2383.62 1.67
c208C15 2 2300.55 3.25 2 2300.55 2300.55 1.09 2 2300.55 2300.55 8.59 2 2300.55 2300.55 1.36
r102C15 5 5412.78 5046.62 5 5412.78 5412.78 1.82 5 5412.78 5412.78 0.09 5 5412.78 5412.78 4.27
r105C15 4 4336.15 5.34 4 4336.15 4336.15 1.03 4 4336.15 4336.15 0.86 4 4336.15 4336.15 2.03
r202C15 2 2361.51 5045.36 2 2358.00 2364.90 22.71 1 1507.32 1677.46 45.16 2 2358.00 2358.00 4.00
r209C15 1 1313.24 3088.27 1 1313.24 1313.24 12.23 1 1313.24 1313.24 10.98 1 1313.24 1313.24 2.83
rc103C15 4 4397.67 245.61 4 4397.67 4397.67 0.24 4 4397.67 4397.67 0.21 4 4397.67 4397.67 2.73
rc108C15 3 3370.25 822.47 3 3370.25 3370.25 41.34 3 3370.25 3370.25 0.24 3 3370.25 3370.25 1.87
rc202C15 2 2394.39 603.76 2 2394.39 2394.39 0.48 2 2394.39 2394.39 19.65 2 2394.39 2394.39 2.72
rc204C15 1 1403.38 5046.60 1 1382.22 1385.72 47.83 1 1382.22 1382.55 50.49 1 1382.22 1384.83 19.75
avg 2.06 2338.38 1331.37 2.06 2337.70 2337.99 8.17 2.03 2314.07 2319.08 7.62 2.06 2337.70 2337.77 2.20

### v2

@@ IV-A3 Discussion
Compared to existing purely sequential insertion procedures [17, 23, 24], PSSI can typically obtain higher-quality feasible routes with reasonable additional computational cost, as demonstrated in the experiments (see Section V-E). The applicability of PSSI extends beyond EVRP-TW-SPD to various EVRP variants sharing certain features: i) partial recharges, ii) customers with time windows, and iii) a homogeneous network of charging stations. For instance, PSSI could be applied to scenarios involving multiple depots, heterogeneous vehicle fleets, or even dynamic routing where customer demands or traffic conditions change in real-time.
@@ IV-B Cross-Domain Neighborhood Search (CDNS)
We propose cross-domain neighborhood search (CDNS), a novel local search procedure that explores two distinct solution spaces simultaneously: the solution space of EVRP-TW-SPD and the solution space of its non-electric counterpart VRP-TW-SPD [21], where some solutions may be infeasible for EVRP-TW-SPD. The non-electric counterpart, VRP-TW-SPD, only excludes the electricity constraint (5g) and, as a result, does not impose any vehicle distance limit. This cross-domain exploration offers two key advantages. First, by expanding beyond the feasible solution space of EVRP-TW-SPD, CDNS potentially discovers higher-quality solutions that might be overlooked by conventional single-domain local search. Second, exploring the solution space of VRP-TW-SPD enables the efficient reuse of well-established local search procedures for VRP-TW-SPD. We will further discuss the broader implications of this cross-domain approach beyond EVRP-TW-SPD at the end of Section IV-B.
@@ 15       $\bm{S}\leftarrow$ find the best-improving solution in the EVRP-TW-SPD neighborhood of $\bm{S}$;
Algorithm 2 outlines the framework of CDNS. It consists of two sub-procedures. The first, called aggressive local search (ALS), explores the solution space of VRP-TW-SPD. As its name suggests, it temporarily ignores the electricity constraint (5g) and repeatedly applies existing local search procedures for VRP-TW-SPD to find the solution with maximum improvement (i.e., the best-improving solution) in the neighborhood of the current solution (lines 22). When ALS can no longer improve the solution, CDNS switches to conservative local search (CLS) which searches for improvements within the solution space of EVRP-TW-SPD (lines 22). Below we detail these two sub-procedures.
@@ IV-B1 Aggressive Local Search (ALS)
When a feasible solution $\bm{S}$ to EVRP-TW-SPD enters ALS (lines 22), the first step is to remove all charging stations in $\bm{S}$, resulting in a feasible solution $\bar{\bm{S}}$ to VRP-TW-SPD (line 2). Then, an existing local search procedure for VRP-TW-SPD is repeatedly applied to find the best-improving solution in the neighborhood of $\bar{\bm{S}}$ (line 2). Here, we employ the local search procedure of MATE [21], one of the state-of-the-art algorithms for solving VRP-TW-SPD. This procedure involves multiple move operators and constant-time move evaluation (see the paper of MATE for details). Then, for each found best-improving solution, PSSI is first applied to transform it into a feasible EVRP-TW-SPD solution (lines 22). If this transformation is not possible, then ALS would terminate (i.e., the second termination condition in line 2) since further exploration of this solution would not lead to a feasible electric version. If successful, the resulting feasible solution would replace the current solution $\bm{S}$ when it has better quality (line 2). ALS would terminate when no further improvement is possible (i.e., the first termination condition in line 2).
@@ IV-B2 Conservative Local Search (CLS)
Unlike ALS, CLS performs local search directly in the solution space of EVRP-TW-SPD. It repeatedly identifies the best-improving solution in the current solution’s neighborhood until no further improvement is possible (lines 22). The neighborhood is defined by the following five move operators. Note that all solutions generated by applying any of the five operators to $\bm{S}$ are evaluated, with the electricity constraint (5g) checked to ensure feasibility.
@@ IV-B3 Discussion
By exploring two distinct solution spaces, CNDS typically finds higher-quality solutions than using either solution space alone, as demonstrated in our experiments (see Section V-E). The idea of CDNS can be readily generalized to other EVRP variants by simultaneously exploring both electric and non-electric neighborhoods. This dual-domain approach is particularly valuable since non-electric variants are typically more extensively studied in the literature, providing a rich pool of well-established local search procedures that can be adapted for solving various EVRPs efficiently.
@@ 6       $\bm{S}^{\prime}\leftarrow\text{CDNS}(\bm{S}^{\prime})$;
9      if $\bm{S}$ is not improved in the last $G_{1}$ iterations then
@@ 21                   $\bm{S}^{\prime}\leftarrow\text{Select\_Best}(\bm{S}^{\prime},\bm{S}_{1},\ldots,\bm{S}_{N})$;
23            until $\bm{S}^{\prime}$ is not improved in the last $G_{2}$ generations;
@@
After initialization, HMA applies the destroy-repair operator as described in [21] (a large-step perturbation) on the individual solution $\bm{S}$ (line 3). This destroy-repair operator involves removing $\rho$ correlated customers from the solution, where the random number $\rho\in[\omega_{1}|\bm{V}|,\omega_{2}|\bm{V}|]$, and $\omega_{1}$, $\omega_{2}$ are parameters with $10<\omega_{1}\leq\omega_{2}<$ 1. The operator then reinserts these customers using a regret-based heuristic that evaluates the expected cost of inserting a node in a future iteration rather than in the current one to assess the quality of potential insertions. The process is followed by CDNS, ensuring that a new local optimum is reached to potentially improve $\bm{S}$ (lines 33). If $\bm{S}$ does not show improvement over a specified number of consecutive iterations, denoted as $G_{1}$, HMA initiates the population-based MS (lines 33).
@@
After initialization, MS enters an evolutionary process (lines 33). In each generation (lines 33), each individual in the population is selected once as parent $\bm{p}_{1}$ and once as parent $\bm{p}_{2}$, in a random order (lines 3 and 3). For each pair of parent solutions, MS generates an offspring using the route-assembly-regret-insertion (RARI) crossover [21] (line 3), and then improves it through CDNS (line 3). After that, the better one among the offspring and $\bm{p}_{1}$ replaces the individual selected as $\bm{p}_{1}$ (line 3). After MS, a new candidate solution $\bm{S}^{\prime}$ is obtained with the aim of improving $\bm{S}$ (lines 33). If the quality of $\bm{S}$ is still not improved by MS, then HMA would output $\bm{S}$ and terminate (line 3); otherwise, HMA proceeds to the next iteration of LNS (line 3).
When presented with a large-scale problem instance (with more than 200 customers) to solve, we make two adjustments to HMA to improve its efficiency. First, only ALS is used in CDNS. Second, problem decomposition is employed in MS. Specifically, the decomposition breaks down the main problem $\bm{G}=(\bm{V}^{\prime},\bm{E})$ into $k$ smaller subproblems $\bm{G}_{1},\ldots,\bm{G}_{k}$ when entering MS (line 3). A subproblem $\bm{G}_{i}=(\bm{V}_{i}^{\prime},\bm{E}_{i})$ with $\bm{V}_{i}^{\prime}=\{0\}\cup\bm{V}_{i}\cup\bm{F}$, $\bm{E}_{i}=\{(p,q)|p,q\in\bm{V}_{i}^{\prime},p\neq q\}$, where $\bm{V}=\bigcup_{i=1}^{k}\bm{V}_{i}\text{ and }\bm{V}_{i}\cap\bm{V}_{j}=\emptyset\ \forall i\neq j,\ 1\leq i,j\leq k$. Here, we use the barycenter clustering decomposition (BCD) [35] to determine the partition of $\bm{V}_{i}$. The barycenter of a route $\bm{R}=(n_{1},n_{2},\ldots,n_{L})$ is defined as $(\overline{x},\overline{y})=\frac{1}{L-1}\sum_{i=1}^{L-1}(x_{i},y_{i}),$ where $(x_{i},y_{i})$ is the Cartesian plane coordinate of node $n_{i}$. For each subproblem $\bm{G}_{i}$, the same MS is applied (lines 33). After all MS procedures are finished, the solutions from all subproblems are assembled to construct a complete solution $\bm{S}^{\prime}$ for the original problem.
@@
Parameter $G_{1}$ $G_{2}$ $N$ $\alpha$ $B$ $sr$ $w_{1},w_{2}$
$akb_{small}$ 20 20 9 3 5 $1.0$ $0.2,0.4$
$akb_{medium}$ 20 20 4 3 5 $0.5$ $0.1,0.2$
$jd$ 20 20 4 3 5 $0.1$ $0.05,0.05$
@@ V Experiments
To evaluate the effectiveness of HMA, we compared it with the recent state-of-the-art algorithms for solving EVRP-TW-SPD, Adapt-CMSA-STD [13] and Adapt-CMSA-SETCOV [19], using the existing benchmark set. Then, we introduced a new benchmark set derived from a real-world application that includes large-scale instances. The evaluation results of Adapt-CMSA-STD, Adapt-CMSA-SETCOV, and HMA on the new benchmark were also reported. Moreover, we used a statistical test to support our claims and measure algorithm robustness. Finally, we conducted a comprehensive ablation study to assess the effectiveness of PSSI, CDNS, and the hybrid search framework. The source code of HMA and the new benchmark set are open-sourced at https://anonymous.4open.science/r/EVRP-TW-SPD-HMA-code-dataset.
@@
Instance name m best time m best avg time m best avg time m best avg time gap
c101C5 2 2257.75 0.43 2 2257.75 2257.75 0.02 2 2257.75 2257.75 0.01 2 2257.75 2257.75 0.38 0.00%
c103C5 1 1175.37 0.41 1 1175.37 1175.37 0.54 1 1175.37 1175.37 0.68 1 1175.37 1175.37 0.21 0.00%
c206C5 1 1242.56 0.58 1 1242.56 1242.56 0.03 1 1242.56 1242.56 0.00 1 1242.56 1242.56 0.19 0.00%
c208C5 1 1158.48 0.08 1 1158.48 1158.48 0.01 1 1158.48 1158.48 0.00 1 1158.48 1158.48 0.24 0.00%
r104C5 2 2136.69 0.02 2 2136.69 2136.69 0.07 2 2136.69 2136.69 0.01 2 2136.69 2136.69 0.93 0.00%
r105C5 2 2156.08 0.03 2 2156.08 2156.08 0.01 2 2156.08 2156.08 0.00 2 2156.08 2156.08 1.28 0.00%
r202C5 1 1128.78 0.06 1 1128.78 1128.78 9.00 1 1128.78 1128.78 0.00 1 1128.78 1128.78 0.24 0.00%
r203C5 1 1179.06 0.03 1 1179.06 1179.06 0.22 1 1179.06 1179.06 0.05 1 1179.06 1179.06 0.17 0.00%
rc105C5 2 2233.77 2.18 2 2233.77 2233.77 0.09 2 2233.77 2233.77 0.04 2 2233.77 2233.77 0.86 0.00%
rc108C5 2 2253.93 0.19 2 2253.93 2253.93 0.01 2 2253.93 2253.93 0.00 2 2253.93 2253.93 0.58 0.00%
rc204C5 1 1176.39 0.25 1 1176.39 1176.39 0.07 1 1176.39 1176.39 0.01 1 1176.39 1176.39 0.21 0.00%
rc208C5 1 1167.98 0.12 1 1167.98 1167.98 0.52 1 1167.98 1167.98 0.03 1 1167.98 1167.98 0.19 0.00%
c101C10 3 3388.25 76.74 3 3388.25 3388.25 0.28 3 3388.25 3388.55 0.33 3 3388.25 3388.25 1.70 0.00%
c104C10 2 2273.93 2.16 2 2273.93 2273.93 0.48 2 2273.93 2273.93 29.02 2 2273.93 2273.93 1.10 0.00%
c202C10 1 1304.06 6.34 1 1304.06 1304.06 0.45 1 1304.06 1304.06 0.10 1 1304.06 1304.06 1.40 0.00%
c205C10 2 2228.28 0.08 2 2228.28 2228.28 10.82 2 2228.28 2228.28 0.03 2 2228.28 2228.28 0.45 0.00%
r102C10 3 3249.19 0.56 3 3249.19 3249.19 18.22 3 3249.19 3249.19 0.01 3 3249.19 3249.19 1.03 0.00%
r103C10 2 2206.12 17.49 2 2206.12 2206.12 5.27 2 2206.12 2206.12 22.01 2 2206.12 2206.12 3.38 0.00%
r201C10 1 1241.51 61.20 1 1241.51 1241.51 22.71 1 1241.51 1241.51 15.97 1 1241.51 1241.51 3.29 0.00%
r203C10 1 1218.21 1.57 1 1218.21 1218.21 14.06 1 1218.21 1218.21 27.17 1 1218.21 1218.21 0.97 0.00%
rc102C10 4 4423.51 0.15 4 4423.51 4423.51 0.24 4 4423.51 4423.51 0.01 4 4423.51 4423.51 3.27 0.00%
rc108C10 3 3345.93 3.72 3 3345.93 3345.93 14.61 3 3345.93 3345.93 0.01 3 3345.93 3345.93 1.29 0.00%
rc201C10 1 1412.86 22756.49 1 1412.86 1412.86 1.97 1 1412.86 1412.86 0.93 1 1412.86 1412.86 6.03 0.00%
rc205C10 2 2325.98 0.13 2 2325.98 2325.98 0.13 2 2325.98 2325.98 0.45 2 2325.98 2325.98 0.74 0.00%
c103C15 3 3348.46 5046.46 3 3348.46 3348.47 48.19 3 3348.46 3348.46 3.76 3 3348.46 3348.46 3.30 0.00%
c106C15 3 3275.13 0.90 3 3275.13 3275.13 4.56 3 3275.13 3275.13 32.34 3 3275.13 3275.13 2.42 0.00%
c202C15 2 2383.62 43.74 2 2383.62 2383.62 12.73 2 2383.62 2393.39 5.27 2 2383.62 2383.62 1.67 0.00%
c208C15 2 2300.55 3.25 2 2300.55 2300.55 1.09 2 2300.55 2300.55 8.59 2 2300.55 2300.55 1.36 0.00%
r102C15 5 5412.78 5046.62 5 5412.78 5412.78 1.82 5 5412.78 5412.78 0.09 5 5412.78 5412.78 4.27 0.00%
r105C15 4 4336.15 5.34 4 4336.15 4336.15 1.03 4 4336.15 4336.15 0.86 4 4336.15 4336.15 2.03 0.00%
r202C15 2 2361.51 5045.36 2 2358.00 2364.90 22.71 1 1507.32 1677.46 45.16 2 2358.00 2358.00 4.00 56.44%
r209C15 1 1313.24 3088.27 1 1313.24 1313.24 12.23 1 1313.24 1313.24 10.98 1 1313.24 1313.24 2.83 0.00%
rc103C15 4 4397.67 245.61 4 4397.67 4397.67 0.24 4 4397.67 4397.67 0.21 4 4397.67 4397.67 2.73 0.00%
rc108C15 3 3370.25 822.47 3 3370.25 3370.25 41.34 3 3370.25 3370.25 0.24 3 3370.25 3370.25 1.87 0.00%
rc202C15 2 2394.39 603.76 2 2394.39 2394.39 0.48 2 2394.39 2394.39 19.65 2 2394.39 2394.39 2.72 0.00%
rc204C15 1 1403.38 5046.60 1 1382.22 1385.72 47.83 1 1382.22 1382.55 50.49 1 1382.22 1384.83 19.75 0.00%
avg 2.06 2338.38 1331.37 2.06 2337.70 2337.99 8.17 2.03 2314.07 2319.08 7.62 2.06 2337.70 2337.77 2.20 1.57%